OpenStudy (anonymous):

I GIVE MEDALS!!!!!!!!! Use inverse operations to solve each equation. Explain each step.

OpenStudy (anonymous):

\[19=\frac{ h }{ 3 }-8\]

OpenStudy (anonymous):

first take the -8 to the other side of the equation make the eq.. 19 + 8 = h/3 27=h/3 as 3 is being divided on one side of the equation transfer it to the other side which means 27 x 3 = h thus h = 81